North Dakota Electric Company estimates its demand trend line (in millions of kilowatt hours) to be: D = 77.0 + 0.45Q,

where Q refers to the sequential quarter number and Q=1 for winter of Year 1. In addition, the multiplicative seasonal factors are as follows:

QuarterFactor (Index)
Winter0.72
Spring1.25
Summer1.45
Fall0.58

In year 26 (quarters 101-104), the energy use for each of the quarters beginning with winter is (round your response to one decimal place):

QuarterEnergy Use
Winter
Spring
Summer
Fall
Answer #1

The demand forecast will be as follows:

D = (77 + 0.45Q) x Multiplicative seasonal factor

Winter = (77+0.45 x 101) x 0.72 = 88.16 million KWH

Spring = (77+0.45 x 102) x 1.25 = 153.63 million KWH

Summer = (77+0.45 x 103) x 1.45 = 178.86 million KWH

Fall = (77+0.45 x 104) x 0.58 = 71.80 million KWH